OpenStudy (anonymous):

on the number line below, which letters show the location for the following numbers: 1/3 Answer: 1 1/3 Answer: A B C I D E F ---------------------------- 0 I I 1 I I I

OpenStudy (anonymous):

1/3=B 1 1/3=D I believe, its kinda hard to tell with what is given

OpenStudy (anonymous):

thanks and sorry im bad at drawing so that was definitely not a choice

OpenStudy (anonymous):

haha no worries

OpenStudy (anonymous):

B=1/3

OpenStudy (anonymous):

D=1 1/3